# Towards Quantifying Long-Range Interactions in Graph Machine Learning: a Large Graph Dataset and a Measurement

### v2

@@ 5 Theoretical Justification of Topologies in City-Networks
In this section, we provide a spectral analysis of over-smoothing in GNNs to justify our dataset design. As such, our main goal is not an in-depth theoretical analysis, since both over-smoothing and spectral analysis of GNNs have been carried out extensively in the literature (Rusch et al., 2023; Wu et al., 2023; Oono & Suzuki, 2020; Rong et al., 2019); instead, our aim is to build on top of these and derive simple results which provide a theoretical grounding of our dataset construction. The section is organized as follows. First, we review the concept of information loss in the limit of infinite GNN layers due to over-smoothing, which effectively vanishes the original node features and makes representations collapse to a value that is only dependent on the graph topology (Section 5.1 and Appendix F.1). Next, we relate the rate at which over-smoothing happens and the eigenvalues of the normalized adjacency operator in linearized GCNs (Section 5.2). Given that eigenvalues with larger absolute values slow down over-smoothing, we analyze how the magnitude of the eigenvalues relates to the graph topology of our datasets. This provides justification that sparse graphs with large diameters promote bigger positive eigenvalues for the normalized adjacency operator and hence are less vulnerable to over-smoothing (Theorem 5 and Appendix F.3). Finally, we argue that less over-smoothing makes it possible for GNNs to capture long-range dependencies in the case of our proposed dataset. All proofs are presented in Appendix F.2.
@@ 5.1 Preliminaries: A Spectral Perspective on Over-smoothing for Linear GCNs
Via linearization (detailed in Appendix F.1), as the number of layers tends to infinity, the graph convolution operation collapses all node features to scalar multiples of $\bm{\tilde{u}}_{N}$, resulting in complete loss of the original feature information, that is, the learned representations suffer from over-smoothing.
@@ 5.2 Over-smoothing Rate on Algebraic Connectivity, Diameter, and Sparsity
We now provide intuition about the types of graph topologies that would mitigate the over-smoothing problem. In particular, sparse graphs—those lacking the small-world effect commonly found in citation networks—and graphs with large diameters tend to experience over-smoothing at a slower rate. We argue that this slower rate of over-smoothing implies a higher likelihood that GNNs can learn long-range dependencies during optimization. We validate this intuition by bounding the eigenvalues of the normalized adjacency matrix using graph properties, then showing how these eigenvalues influence over-smoothing, and thus how controlling these graph properties can reduce the likelihood of over-smoothing. More concretely, we extend (Wu et al., 2019, Theorem 1), which only covers $\lambda_{1}$ and $\lambda_{N}$, to Proposition 4 which we prove in Appendix F.2, covering the normalized algebraic connectivity $\lambda_{N-1}$ and showing that $\lambda_{N-1}\left(\bm{S}_{adj}\right)<\lambda_{N-1}(\tilde{\bm{S}}_{adj})$.
@@
From Table 3, we see that the lower bound in Equation (5) is generally higher for our datasets; this is because the bound is decreasing in $d_{max}$ and increasing in $diam(\mathcal{G})$, showing graphs with large diameter and low maximum degree will be more resilient to over-smoothing. We further provide the rationale behind this lower bound in Appendix F.3.
@@ 6 Theoretical Justification of Jacobian-Based Measurement
Finally, we justify the proposed measurement for analyzing long-range interactions in GNNs, particularly by comparing the mean and the sum of influence scores of neighboring nodes as candidates to represent the total influence score for a focal node $v$, $T_{h}(v)$. Recall that $T_{h}(v)=I_{\text{sum}}(v,h)=\sum_{u:\rho(v,u)=h}I(v,u)$ as in Equation (2). We first introduce definitions for the standard lattice in Euclidean space, $h$-hop shells, and quasi-isometric graphs, and then provide a mathematical analysis of influence score dilution. Due to space limits, we provide the definitions in Appendix H.
@@
The analysis above formally justifies the choice in Equation (2), which considers the aggregate influence of neighboring nodes $T_{h}(v)=I_{\text{sum}}$ as a more reliable measure than the mean $I_{\text{mean}}$, which is susceptible to dilution, particularly in the case of distant neighbors. This is key to the computation of the average total influence in Equation (2) and the influence-weighted receptive field in Equation (3). Finally, note the following:
@@
Proofs are provided in Appendix I, and a more detailed justification of Jacobian-based influence score and connections to existing literature, along with its limitations, can be found in Appendix G.
@@ 7 Conclusion
Dataset and Code Availability
Our dataset, CityNetwork, is now publicly available in the latest version of PyTorch Geometric (PyG) as an InMemoryDataset under torch_geometric.datasets. The Total Influence metric introduced in Section 4 is also released under torch_geometric.utils. We open-sourced the code for dataset construction and benchmarking at https://github.com/LeonResearch/City-Networks, and ensure that the experimental results and figures in this work are reproducible with the provided code.
Appendix A Network Statistics
@@
Appendix B Dataset Details
B.1 Raw Feature Processing
@@
B.2 Long-range Node Labels based on Local Eccentricity
@@
B.3 The Roles of Graph Topology and Spatial Features
@@
Problem of uninformative graph structures on labels.
One catastrophic problem we encountered during the experiment is that, when using the exact eccentricity or a relatively larger $k$ for node labels, the signal becomes nearly independent of graph structures. As illustrated in Figure 6, we can observe a clear correlation between the node labels and their spatial coordinates, since nodes around the city center would generally have lower eccentricity values due to the 2D grid-like topology. This means that the spatial coordinates alone are sufficient for modelling $\varepsilon(v)$, hence weakening the importance of graph structure during modeling. Similarly, some other labeling approaches would also lead to the same problem, such as using a single node as an anchor, then regarding its shortest path distance to other nodes as the node label.
@@ B.4 Choice of the Long-range Level K
Appendix C Experiment Details
@@
C.1 Baselines and Datasets for Comparison
@@
We consider four common GNNs: GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCNII (Chen et al., 2020) for GNNs; and three Graph Transformers: GraphGPS (Rampášek et al., 2022), Exphormer (Shirzad et al., 2023), and SGFormer (Wu et al., 2023) as the baseline models in our experiments.
@@ C.2 Experimental Setups
We run each model for $30k$ maximum epochs using AdamW (Loshchilov & Hutter, 2019) optimizer for sufficient training, during which we record the validation accuracy every $100$ epochs and save the model at the best validation epoch for final testing. Meanwhile, we closely follow the latest GNN tuning technique from Luo et al. (2024), and our hyperparameter search space is summarized in Table 4. Note that for GTs, we do not apply positional encoding due to its impractical computation on our large graphs (which is one of the challenges of applying GTs on large graphs), and use GCN as their internal MPNNs.
@@ C.3 Ablation Studies
To further support our main findings, we test two classical GNN baselines: GCN (Kipf & Welling, 2017) and GraphSAGE (Hamilton et al., 2017); and two GT baselines: Exphormer (Shirzad et al., 2023) and SGFormer (Wu et al., 2023) under different hidden_size=[16, 32, 64, 128] on Paris, Shanghai, and L.A.. The results are presented in Figure 9, where we can observe a consistent pattern with our main findings in Figure 3—the baselines’ performance improves substantially when increasing the number of layers from $2$ to $16$.
@@
C.4 Summary of Baseline Results
For reference, we summarize the best baseline results on our City-Networks in Table 5. Note that due to the global attention module in GraphGPS, it is highly memory-consuming and fails to fit into our GPU memory (4848GB), even with hidden_size=$16$ under num_layers=$16$.
Table 5: Baseline results on City-Networks.
Baseline Type Paris Shanghai Los Angeles London
MPNNs MLP $25.5\pm 0.4$ $28.4\pm 0.6$ $24.1\pm 0.5$ $27.9\pm 0.1$
ChebNet $54.1\pm 0.2$ $66.5\pm 0.1$ $61.4\pm 0.4$ $54.7\pm 0.2$
GCN $53.2\pm 0.3$ $62.1\pm 0.2$ $58.3\pm 0.3$ $50.1\pm 0.7$
GraphSAGE $54.6\pm 0.2$ $68.3\pm 0.5$ $61.4\pm 0.3$ $55.4\pm 0.2$
GAT $51.1\pm 0.3$ $68.0\pm 0.5$ $59.5\pm 0.3$ $52.0\pm 0.3$
GCNII $51.3\pm 0.2$ $61.5\pm 0.4$ $56.0\pm 0.3$ $48.2\pm 0.3$
DropEdge $48.2\pm 0.2$ $60.8\pm 0.4$ $55.5\pm 0.3$ $45.0\pm 0.3$
GTs GraphGPS $52.1\pm 0.6$ $63.0\pm 0.5$ $59.8\pm 0.5$ OOM
Exphormer $55.1\pm 0.8$ $70.2\pm 0.4$ $63.8\pm 0.6$ $49.5\pm 0.4$
SGFormer $52.0\pm 0.8$ $64.1\pm 0.3$ $60.1\pm 0.7$ $48.3\pm 0.3$
Appendix D Computational Complexity in R

### v3

@@ 5 Theoretical Justification of Topologies in City-Networks
In this section, we provide a spectral analysis of over-smoothing in GNNs to justify our dataset design. As such, our main goal is not an in-depth theoretical analysis, since both over-smoothing and spectral analysis of GNNs have been carried out extensively in the literature (Rusch et al., 2023; Wu et al., 2023; Oono and Suzuki, 2020; Rong et al., 2019); instead, our aim is to build on top of these and derive simple results which provide a theoretical grounding of our dataset construction. The section is organized as follows. First, we review the concept of information loss in the limit of infinite GNN layers due to over-smoothing, which effectively vanishes the original node features and makes representations collapse to a value that is only dependent on the graph topology (Section 5.1 and Appendix G.1). Next, we relate the rate at which over-smoothing happens and the eigenvalues of the normalized adjacency operator in linearized GCNs (Section 5.2). Given that eigenvalues with larger absolute values slow down over-smoothing, we analyze how the magnitude of the eigenvalues relates to the graph topology of our datasets. This provides justification that graphs with large diameters promote bigger positive eigenvalues for the normalized adjacency operator and hence are less vulnerable to over-smoothing (Theorem 5 and Appendix G.3). Finally, we argue that less over-smoothing makes it possible for GNNs to capture long-range dependencies in the case of our proposed dataset. All proofs are presented in Appendix G.2.
@@ 5.1 Preliminaries: A Spectral Perspective on Over-smoothing for Linear GCNs
Via linearization (detailed in Appendix G.1), as the number of layers tends to infinity, the graph convolution operation collapses all node features to scalar multiples of $\bm{\tilde{u}}_{N}$, resulting in complete loss of the original feature information, that is, the learned representations suffer from over-smoothing.
@@ 5.2 Over-smoothing Rate on Algebraic Connectivity, Diameter, and Sparsity
We now provide intuition about the types of graph topologies that would mitigate the over-smoothing problem. In particular, sparse graphs—those lacking the small-world effect commonly found in citation networks—and graphs with large diameters tend to experience over-smoothing at a slower rate. We argue that this slower rate of over-smoothing implies a higher likelihood that GNNs can learn long-range dependencies during optimization. We validate this intuition by bounding the eigenvalues of the normalized adjacency matrix using graph properties, then showing how these eigenvalues influence over-smoothing, and thus how controlling these graph properties can reduce the likelihood of over-smoothing. More concretely, we extend (Wu et al., 2019, Theorem 1), which only covers $\lambda_{1}$ and $\lambda_{N}$, to Proposition 4 which we prove in Appendix G.2, covering the normalized algebraic connectivity $\lambda_{N-1}$ and showing that $\lambda_{N-1}\left(\bm{S}_{adj}\right)<\lambda_{N-1}(\tilde{\bm{S}}_{adj})$.
@@
From Table 4, we see that the lower bound in equation 5 is generally higher for our datasets; this is because the bound is decreasing in $d_{max}$ and increasing in $diam(\mathcal{G})$, showing graphs with large diameter and low maximum degree will be more resilient to over-smoothing. We further provide the rationale behind this lower bound in Appendix G.3, as well as its relations to the Braess paradox and previous study (Jamadandi et al., 2024) regarding over-smoothing and graph sparsity in Appendix G.4.
@@ 6 Theoretical Justification of Jacobian-Based Measurement
Finally, we justify the proposed measurement for analyzing long-range interactions in GNNs, particularly by comparing the mean and the sum of influence scores of neighboring nodes as candidates to represent the total influence score for a focal node $v$, $T_{h}(v)$. Recall that $T_{h}(v)=I_{\text{sum}}(v,h)=\sum_{u:\rho(v,u)=h}I(v,u)$ as in equation 2. We first introduce definitions for the standard lattice in Euclidean space, $h$-hop shells, and quasi-isometric graphs, and then provide a mathematical analysis of influence score dilution. Due to space limits, we provide the definitions in Appendix I.
@@
The analysis above formally justifies the choice in equation 2, which considers the aggregate influence of neighboring nodes $T_{h}(v)=I_{\text{sum}}$ as a more reliable measure than the mean $I_{\text{mean}}$, which is susceptible to dilution, particularly in the case of distant neighbors. This is key to the computation of the average total influence in equation 2 and the influence-weighted receptive field in equation 3. Finally, note the following:
@@
Proofs are provided in Appendix J, and a more detailed justification of Jacobian-based influence score and connections to existing literature, along with its limitations, can be found in Appendix H.
@@ 7 Conclusion
Acknowledgment
Huidong Liang is funded by the ESRC Grand Union Doctoral Training Partnership and the Oxford-Man Institute of Quantitative Finance.
Appendix A Dataset and Code Availability
Our dataset, CityNetwork, and our long-range measurement, total_influence, are both publicly available on PyTorch Geometric (PyG 2.7.0) under torch_geometric.datasets and torch_geometric.utils. We also open-sourced the code for dataset construction and benchmarking at https://github.com/LeonResearch/City-Networks.
Appendix B Network Statistics
@@
Appendix C Dataset Details
C.1 Raw Feature Processing
@@
C.2 Long-range Node Labels based on Local Eccentricity
@@
Figure 6: Results on Paris where labels are derived with travel time being the edge weight.
Discussions on more realistic estimations of accessibility.
While our approach gives us a clean and controllable setup to study long-range topological interactions, it inevitably neglects some of the complexities inherent in real-world transportation dynamics, in which other factors like road speed, capacity, and traffic congestion strongly influence how “accessible” one area is from another. However, we believe our framework can be naturally extended by using road length and speed limit to approximate travel time when defining the edge weight in the eccentricity from equation 1. Here we present some preliminary results on Paris under this strategy. The results in Figure 6 suggest a similar rising trend in performance as we increase the model’s depth, which is consistent with our statements in the main text. We note that other travel-time proxies, such as data from the Google Maps API, may provide a closer approximation to real-world conditions, and we leave such directions for future work.
@@
C.3 The Roles of Graph Topology and Spatial Features
@@
Problem of uninformative graph structures on labels.
One catastrophic problem we encountered during the experiment is that, when using the exact eccentricity or a relatively larger $k$ for node labels, the signal becomes nearly independent of graph structures. As illustrated in Figure 7, we can observe a clear correlation between the node labels and their spatial coordinates, since nodes around the city center would generally have lower eccentricity values due to the 2D grid-like topology. This means that the spatial coordinates alone are sufficient for modeling $\varepsilon(v)$, hence weakening the importance of graph structure for the prediction task. Similarly, some other labeling approaches would also lead to the same problem, such as using a single node as an anchor, then regarding its shortest path distance to other nodes as the node label.
Empirical validation.
In Section 3, we use an MLP on all four city networks that only uses node features such as geographical coordinates, land use, etc. The result in Figure 3 shows a significant performance gap between MLP and other graph models, indicating that using spatial information alone is insufficient for our task. To further show the sensitivity of the baseline models to geographic coordinates, we test two GNNs (GCN and GraphSAGE) and two GTs (Exphormer and SGFormer) on Paris and Shanghai, with coordinates masked in node features. At the same time, we also test MLP on these two cities with coordinates only, and summarize the results in Table 5 below. Compared to the original results with all spatial features, we can observe a slight performance drop across both GNNs and GTs after removing the geographical coordinates, while for MLP, the results indicate that geographical coordinates alone are not sufficient for modelling our long-range signal.
Table 5: Baseline results on Paris and Shanghai with all features vs. coordinates only (MLP) and coordinates masked (GNNs and GTs).
Paris Shanghai
Model all features coords. only / masked all features coords. only / masked
MLP 25.5 $\pm$ 0.4 12.5 $\pm$ 0.5 28.4 $\pm$ 0.6 15.2 $\pm$ 0.7
GCN 53.2 $\pm$ 0.3 51.4 $\pm$ 0.4 62.1 $\pm$ 0.2 61.3 $\pm$ 0.4
GraphSAGE 54.6 $\pm$ 0.2 52.3 $\pm$ 0.3 68.3 $\pm$ 0.5 66.5 $\pm$ 0.4
Exphormer 55.1 $\pm$ 0.8 53.5 $\pm$ 0.4 70.2 $\pm$ 0.4 67.4 $\pm$ 0.5
SGFormer 52.0 $\pm$ 0.8 51.3 $\pm$ 0.7 64.1 $\pm$ 0.3 62.8 $\pm$ 0.4
Discussion on LRGB.
@@
C.4 Choice of the Long-range Level K
@@
Appendix D Experimental Details
@@
D.1 Baselines and Datasets for Comparison
@@
We consider four common GNNs: GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCNII (Chen et al., 2020) for GNNs; and three Graph Transformers: GraphGPS (Rampášek et al., 2022), Exphormer (Shirzad et al., 2023), and SGFormer (Wu et al., 2023) as the baseline models in our experiments.
@@ D.2 Experimental Setups
We run each model for $30k$ maximum epochs using AdamW (Loshchilov and Hutter, 2019) optimizer for sufficient training, during which we record the validation accuracy every $100$ epochs and save the model at the best validation epoch for final testing. Meanwhile, we closely follow the latest GNN tuning technique from Luo et al. (2024), and our hyperparameter search space is summarized in Table 6. Note that for GTs, we do not apply positional encoding due to its impractical computation on our large graphs (which is one of the challenges of applying GTs on large graphs), and use GCN as their internal MPNNs.
@@ D.3 Ablation Studies
To further support our main findings, we test two classical GNN baselines: GCN (Kipf and Welling, 2017) and GraphSAGE (Hamilton et al., 2017); and two GT baselines: Exphormer (Shirzad et al., 2023) and SGFormer (Wu et al., 2023) under different hidden_size=[16, 32, 64, 128] on Paris, Shanghai, and L.A.. The results are presented in Figure 10, where we can observe a consistent pattern with our main findings in Figure 3—the baselines’ performance improves substantially when increasing the number of layers from $2$ to $16$.
@@
Figure 12: Results for models at deeper depths $L\in[32,48,64]$ beyond the ground-truth range of $16$ hops. For illustration purposes, we also show results at model depths of $[2,4,8,16]$.
@@
Deeper depths beyond the ground-truth range.
While adopting a fixed ground-truth range prevents our long-range task from falling into an unbounded "the longer the communication path, the better the model performance" setting, we additionally benchmark GNNs with depths much deeper than the ground-truth task range (16th hop) here for a more complete analysis of the model’s behavior. In particular, we test GCN and GraphSAGE with depth=[32, 48, 64] on Paris, Shanghai, and L.A. under the same experimental settings (GTs are not tested here due to OOM at these depths), where results are presented in Figure 12. As expected, we can observe that both GNNs, even with residual connections and batch norm enabled, start to suffer from over-smoothing when model depth exceeds the ground-truth range of 16.
In addition, we also show the results of our per-hop influence measurement and $R$ in Figure 13 and Table 7, respectively, where we compare the behaviors of these two deep GNNs on our city networks to those on the other common graph datasets. The results reveal a consistent trend with our findings in Section 4 that the influence scores decay at a much slower rate on our city networks compared to those on the existing graph datasets in the literature. It is also worth noting that the current model depth of 64 exceeds the diameters of those common graph benchmarks, which leads to the “cut-off” pattern in Figure 12.
Figure 13: Normalized average total influence $\bar{T}_{h}/\bar{T}_{0}$ averaged across nodes, where the underlying models are trained with 64 layers—much deeper than the ground-truth range of the task.
Table 7: Average size of the influence-weighted receptive field $R$ across datasets on GCN and GraphSAGE, where both models are trained with 64 layers with residual connections and batch norm.
Model Paris Shanghai L.A. Cora Amazon PascalVOC
GCN 8.61 10.21 9.64 5.17 6.51 7.06
GraphSAGE 8.34 10.72 10.54 3.74 2.78 7.62
D.4 Results on RingTransfer
The RingTransfer (Bodnar et al., 2021) experiment is used for testing long-range dependency in GNNs under inductive settings with small ring-like graphs. In this section, we will analyze the behaviors of different graph models on this synthetic task using our influence metric.
Figure 14: Influence scores on RingTransfer.
Task description.
Each graph in the RingTransfer dataset is a ring of $N$ nodes with only two nodes marked: the source node and the target node, which are placed at opposite sides of the ring (i.e., at a distance of the diameter $N/2$). All nodes on the ring will have a constant feature vector except for the source node, which has a one-hot encoding of its label. The task is to train a model such that the target node’s representation predicts the source’s label, which requires the model to propagate long-range information from the source node to the target node.
Setups.
We followed the same setups by Bodnar et al. (2021) and adopted k=16 (i.e., a task range of 8 hops) with a model depth of 8 in our setting, as we observed a similar phenomenon in their paper that GNNs start to deteriorate after this point. Meanwhile, since the diameter of rings matches the true task range of $N/2$ in RingTransfer, there will be no gradient w.r.t. nodes from hops > $N/2$. Therefore, we test models at depth up to the ground-truth range in this setting. In particular, we evaluate GCN, GraphSAGE, and Exphormer with training/validation/test splits of $5k/1k/1k$ graphs, where all models can achieve $100\%$ accuracy on the testing set. We then apply our measurement on the target node only for each graph, and then report the average over the testing set.
Results and discussion.
The results for $R$ and per-hop influence are summarized in Figure 14. As expected, we can observe a strong influence at hop 8 for GCN and GraphSAGE; while for Exphormer, as it adopts global virtual node and expander graph operations in its attention mechanism, the source node can be effectively reached within a single hop. Therefore, we observe a higher influence on the first few hops compared to more distant hops since the underlying graph has been modified. In addition, we have also tested models with deeper depth of 16 layers, where the result shows a similar pattern to that in Figure 14, except for the 0 influence after hop 8 (as explained above).
While we do observe different influence patterns due to different model designs, we’d also like to point out that this is because RingTransfer is too simplistic and relies solely on long-range interactions, such that simple operations (e.g., rewiring, virtual node, etc.) will convert it into a short-range task.
D.5 More Results on Heterophilic Datasets
As pointed out by Arnaiz-Rodriguez and Errica (2025), the heterophilic datasets are often misused in the graph machine learning literature to evaluate long-range interaction (Arnaiz-Rodríguez et al., 2022; Maskey et al., 2023; Giraldo et al., 2023; Singh et al., 2025). However, there is no empirical study that compares the range of tasks between homophilic and heterophilic graphs.
In Section 3 and 4, we have tested baselines and quantified their long-range influence on a representative heterophilic dataset, Amazon-Ratings. To further support our findings that heterophily does not empirically imply long-rangeness, we proceed to show more results on the following heterophilic datasets that are commonly used in the literature: Roman-Empire (Platonov et al., 2023), Wisconsin, Texas, and Cornell (Pei et al., 2020). The results for per-hop influence and $R$ are presented in Figure 15 and Table 8, respectively, where we can observe from these two diagnostics that heterophilic datasets generally exhibit a short-range pattern compared to our city networks across different models, which is consistent with the findings in our paper.
Figure 15: Normalized average total influence $\bar{T}_{h}/\bar{T}_{0}$ averaged across nodes at different hops on heterophilic datasets, where results on other datasets are also presented for comparison purposes.
Table 8: Average size of the influence-weighted receptive field $R$ across heterophilic datasets and models with 16 layers, where results on other datasets are also presented for comparison purposes.
Model L.A. Cora arxiv PascalVOC Amazon Roman Wisconsin Texas Cornell
GCN 5.36 2.56 1.34 3.38 1.92 2.04 1.28 1.11 1.54
GraphSAGE 5.44 2.37 1.40 2.99 1.80 1.89 0.97 0.89 0.96
GraphGPS 7.92 2.65 OOM 7.14 6.86 6.37 1.03 0.77 0.98
SGFormer 4.03 3.21 1.21 NA 2.46 2.33 0.14 0.24 0.17
Note that Wisconsin, Texas, and Cornell are small heterophilic graphs (around 200 nodes) and all have a small diameter of 8, which explains their extreme short-range behaviors compared to other heterophilic datasets like Amazon-Ratings and Roman-Empire.
Appendix E Computational Complexity